Shows the car, which takes a horizontal turn with speed V = 36 km / h, less than the allowed speed. Calculate the frictional force of rest, which is exerted on each wheel of the car, when it is known that the turning radius is R = 50 m. The mass of the machine is m = 1.2t.


1
Expert's answer
2021-01-19T07:10:24-0500

"36(km\/h)=10(m\/s)"


"4F_f=m\frac{v^2}{r}\to F_f=\frac{mv^2}{4r}=\frac{1200\cdot10^2}{4\cdot50}=600(N)"